OpenStudy (anonymous):

3.6% of 50 is what ? 1.8 yes , no ?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

ok your not just telling me right?

OpenStudy (anonymous):

3.6% = 3.6/100 so (50)(3.6/100) = 3.6/2 = 1.8

OpenStudy (anonymous):

ok great thanks

OpenStudy (anonymous):

your welcome

zepdrix (zepdrix):

yay suzy \c:/